OpenStudy (anonymous):

4x-10y=-3 12x+5y=12 Solve for X? can someone explain

OpenStudy (jh3power):

316x=15y y=316/15 x 4x-10y=4x-10(316/15)=4x-632/3=12 4x=12+632/3 x=3+158/3

OpenStudy (jh3power):

i'm stuck..it shouldn't be this complicated.. i must've made a mistake.

OpenStudy (anonymous):

thats exactly how to question was given to me.

OpenStudy (anonymous):

the*

OpenStudy (jh3power):

but there's nothing wrong with my solution.. the question must be wrong for somewhat reason..

OpenStudy (jh3power):

there happens to be such circumstances, really.

OpenStudy (jh3power):

oh wait.. mistake!

OpenStudy (anonymous):

do you want to see the answers the gave me?

OpenStudy (jh3power):

4x-10(316/15 x)=4x-632/3 x=12 -620/3 x=12 x=-36/620 x=-9/154

OpenStudy (jh3power):

yep, sure

OpenStudy (anonymous):

A) -15/28 B)-9/7 C) 3/4 D)3/5

OpenStudy (jh3power):

I guarantee that the question itself is actually wrong, lol..

OpenStudy (anonymous):

lol